Contingency Tables

Instead of Venn Diagrams, we can use contingency tables to show probabilities of combinations of events:

Pasted image 20240924102149.png

Here the upper-left entry is P(A∩B), and similarly for the other cells. Since the cells all union to S, their probabilities must add to 1. The sum of the A row is P(A) and the sum of the B column is P(B).

Example

Using the residential properties example, we have P(I),P(T),P(I∩T) giving:

| | T | T′ | |
| ---- | --- | ---- | --- |
| I | .5 | | .6 |
| I′ | | | |
| | .8 | | |

Since P(T′)=1−P(T)=.2 and P(I′)=1−P(I)=0.4 then we know the other totals. Doing similar deductions with the other properties of probabilty will give the total table:

| | T | T′ | |
| ---- | --- | ---- | --- |
| I | .5 | .1 | .6 |
| I′ | .3 | .1 | .4 |
| | .8 | .2 | 1.0 |

If you wanted to use the table to get say P(I∪T) then you would just use the addition rule:

P(I∪T)=P(I)+P(T)−P(I∩T)=.6+.8−.5=.9

and so on.

References

  1. [[Matthew A. Carlton, Jay L. Devore - Probability with STEM Applications-Wiley (2020).pdf#page=48]]